OpenStudy (anonymous):

how do you correctly simplifies the expression: (2^2 * 6^0 / 7)^3 ???

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

can you help @KamiBug

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@tester97

OpenStudy (alexandervonhumboldt2):

2^(2*3)*6^(0/7*3)=2^6*1=64

Directrix (directrix):

(2^2 * 6^0 / 7)^3